# Nambu-Goto equation from three-dimensional gravity

### v1

@@ Abstract
We demonstrate that the solutions of three-dimensional gravity obtained by gluing two copies of a spacetime across a junction constituted of a tensile string are in one-to-one correspondence with the solutions of the Nambu-Goto equation in the same spacetime up to a finite number of rigid deformations. The non-linear Nambu-Goto equation satisfied by the average of the embedding coordinates of the junction emerges directly from the junction conditions along with the rigid deformations and corrections due to the tension. Therefore, the equivalence principle generalizes non-trivially to the string. Our results are valid both in three-dimensional flat and AdS spacetimes. In the context of AdS3/CFT2 correspondence, our setup could be used to describe a class of interfaces in the conformal field theory featuring relative time reparametrization at the interface which encodes the solution of the Nambu-Goto equation corresponding to the bulk junction.
@@ 1 Introduction
We examine this question in pure three-dimensional gravity in flat and AdS spaces. We consider solutions which are obtained by gluing two copies of a spacetime $\mathcal{M}$ across a junction constituted of a string with a finite tension. We find that such solutions of pure gravity are in one-to-one correspondence with the solutions of the Nambu-Goto equation for the motion of the string in $\mathcal{M}$ up to corrections due to the finite tension and a finite number of possible rigid deformations. By rigid deformations, we mean additional contributions that arise due to the worldsheet isometries or are related to displacements of a hypersurface which preserve its extrinsic curvature. In particular, we show that the Nambu-Goto equation is recovered in the tensionless limit in absence of the rigid deformations.
@@ 2 Setup and statement of results
Consider two copies of a $2+1$-dimensional Einstein manifold $\mathcal{M}$, which is locally flat or an AdS space. We label the copies of $\mathcal{M}$ as $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$. Let $\Sigma_{1}$ and $\Sigma_{2}$ be two $1+1$-dimensional hypersurfaces embedded in $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$, respectively, and splitting each of these spacetimes into two halves. We will study solutions of Einstein’s equations obtained by gluing one of the halves of $\mathcal{M}_{1}$ with one of the halves of $\mathcal{M}_{2}$ by identifying points in $\Sigma_{1}$ and $\Sigma_{2}$ such that the induced metrics and extrinsic curvatures satisfy the Israel junction conditions Israel:1966rt . The methodology is similar to that adopted in Bachas:2020yxv , except that we will allow more general boundary conditions and also not restrict ourselves to linear perturbations about a simple solution. See Fig. 1 for an illustration of the setup in AdS.
@@
$t_{1}(\tau,\sigma)=\tau-\tau_{a}(\tau,\sigma),\quad z_{1}(\tau,\sigma)=\sigma-\sigma_{a}(\tau,\sigma),\quad x_{1}=f_{1}\left(t_{1}(\tau,\sigma),x_{1}(\tau,\sigma)\right),$
$t_{2}(\tau,\sigma)=\tau+\tau_{a}(\tau,\sigma),\quad z_{2}(\tau,\sigma)=\sigma+\sigma_{a}(\tau,\sigma),\quad x_{2}=f_{2}\left(t_{2}(\tau,\sigma),x_{2}(\tau,\sigma)\right).$ (4)
@@
$x_{s}(\tau,\sigma)=\frac{1}{2}\left(f_{1}\left(t_{1}(\tau,\sigma),x_{1}(\tau,\sigma)\right)+f_{2}\left(t_{2}(\tau,\sigma),x_{2}(\tau,\sigma)\right)\right),$
$x_{a}(\tau,\sigma)=\frac{1}{2}\left(f_{2}\left(t_{2}(\tau,\sigma),x_{2}(\tau,\sigma)\right)-f_{1}\left(t_{1}(\tau,\sigma),x_{1}(\tau,\sigma)\right)\right),$ (5)
@@
Henceforth, we will refer to $\lambda$ as the tension. The two junction conditions give six equations. However, the Brown-York tensor tensor is identically conserved on any hypersurface of an Einstein manifold by virtue of the Gauss-Codazzi equations and this guarantees that the traceless part of the Brown-York tensor is continuous at the junction even when $\lambda\neq 0$. Therefore, the gluing conditions (7) give only one independent equation which together with (6) amount to four equations for determining the four variables, namely $\tau_{a}$, $\sigma_{a}$, $f_{1}$ and $f_{2}$.
We will see that a solution of the junction conditions depends on the tension $\lambda$ and a finite number of integration constants $\delta_{i}$, that we call rigid deformations. The latter are associated with worldsheet isometries and rigid displacements of the hypersurface which do not change its extrinsic curvature.
@@ 3.1 Solving the junction conditions perturbatively
It is convenient to use the average and relative coordinates $x_{s}$ and $x_{a}$ as defined in (2). Obviously, the coefficients of their perturbative expansions are
@@
where $\dot{\phantom{a}}$ and $\phantom{a}{}^{\prime}$ imply partial derivaties w.r.t. $\tau$ and $\sigma$, respectively. At this order, the junction conditions (7) give
@@
It is easy to see that $\alpha_{k}$, $\beta_{k}$ and $\gamma_{k}$ are rigid displacement parameters that preserve the extrinsic curvature of a hypersurface. It will turn out that the higher orders in the expansion will not introduce new parameters, so that in total $\alpha_{h}$, $\beta_{h}$, $\gamma_{h}$, $\alpha_{k}$, $\beta_{k}$ and $\gamma_{k}$ give six rigid parameters which together with the initial conditions for $x_{s}$ uniquely specify a solution of the junction conditions to all orders. By assumption, all these six rigid parameters are small, i.e. $\mathcal{O}(\varepsilon)$.
@@ 4.1 Solving the junction conditions perturbatively
One crucial point is that we do not impose any boundary conditions on $\tau_{a}$ and $\sigma_{a}$. In fact, these will be given by two rigid parameters related to worldsheet isometries at the leading order. [^3] This is in contrast to the setup of Bachas:2020yxv , where they imposed Dirichlet boundary conditions on all the variables and found a unique solution for each given frequency in pure AdS. The diffeomorphisms related to these two parameters preserve the worldsheet boundary at $\sigma=\infty$ but induce relative time-reparametrization at the junction since $\tau_{a}$ does not vanish at the boundary of AdS. Unlike the case of flat space, our perturbative expansion will be valid near the boundary although it will be unreliable near the horizon.
@@
Just like in the case of flat space, the parameters $\alpha_{k}$, $\beta_{k}$, and $\gamma_{k}$ parametrize the rigid infinitesimal deformations of a hypersurface which preserves its extrinsic curvature. However, imposing the Dirichlet boundary condition on both $x_{1}$ and $x_{2}$, leads to the boundary conditions (54) for $x_{a,1}$ implying that it should vanish at $\sigma\rightarrow\infty$. This can be satisfied if
@@ 5 Discussion
Our general solutions which correspond to the Nambu-Goto solutions up to the rigid deformations should be interpreted as general state-dependent interfaces in the dual conformal field theory. Note that the Nambu-Goto equation has no intrinsic dimensionful parameters, and the quasinormal mode spectrum that characterize the solutions is determined by the mass of the background spacetime, i.e. the dual state. When we impose the Dirichlet boundary condition identically on both the hypersurfaces (i.e. for both $x_{s}$ and $x_{a}$), the solutions should holographically represent dynamical interfaces between two copies of the same state in the dual CFT, particularly two thermal states with the same temperature since the BTZ black hole is dual to a thermal state. No transfer of energy and momentum occur through the interfaces. Nevertheless, the interfaces have their own dynamics, which are characterized by the solutions of the Nambu-Goto equations and the rigid parameters of the bulk solution. All these parameters and the solutions of the Nambu-Goto equation together with the tension determine the correlation functions of the full system. In this aspect, it could be important to recall that two sides of the interface necessarily have a relative time-reparametrization (given by the asymptotic limit of $\tau_{a}$) which can decode the non-trivial Nambu-Goto solution in the bulk. This feature is not present in a usual defect. In this context, it is important to understand such interphases in the vacuum which correspond to bulk junctions in pure AdS3.
@@
At the classical level, our constructions extend beyond three spacetime dimensions. As for instance, we can consider a $2+1$-dimensional junction in $3+1$-dimensional flat space in which one spatial direction is compactified to a circle and in which one spatial direction of the junction wraps the circle. Dimensional reduction can give a three-dimensional setup studied here but with a massless scalar field and an $U(1)$ gauge field added in the bulk. The latter should induce additional dynamics on the worldsheet. Similarly, Geroch’s reduction method adapted to Einstein spacetimes Leigh:2014dja can be possibly generalized to dimensionally reduce the $3+1$-dimensional spacetimes with junctions to $2+1$-dimensions where the dynamics is described by gravity coupled to a non-linear sigma model in which the target space is $R\times H_{2}$. [^5] In fact, such branes could lead to stable four-dimensional de-Sitter spaces as discussed recently in Ghosh:2018fbx . Additional dynamics should also be induced on the worldsheet in this case. It will be interesting to see if such worldsheet theories are amenable to quantization.

### v2

@@ Abstract
We demonstrate that the solutions of three-dimensional gravity obtained by gluing two copies of a spacetime across a junction constituted of a tensile string are in one-to-one correspondence with the solutions of the Nambu-Goto equation in the same spacetime up to a finite number of rigid deformations related to worldsheet and spacetime isometries. The non-linear Nambu-Goto equation satisfied by the average of the embedding coordinates of the junction emerges directly from the junction conditions along with the rigid deformations and corrections due to the tension. Therefore, the equivalence principle generalizes non-trivially to the string. Our results are valid both in three-dimensional flat and AdS spacetimes. In the context of AdS3/CFT2 correspondence, our setup could be used to describe a class of interfaces in the conformal field theory featuring relative time reparametrization at the interface which encodes the solution of the Nambu-Goto equation corresponding to the bulk junction.
@@ 1 Introduction
We examine this question in pure three-dimensional gravity in flat and AdS spaces. We consider solutions which are obtained by gluing two copies of a spacetime $\mathcal{M}$ across a junction constituted of a string with a finite tension. We find that such solutions of pure gravity are in one-to-one correspondence with the solutions of the Nambu-Goto equation for the motion of the string in $\mathcal{M}$ up to corrections due to the finite tension and a finite number of possible rigid deformations. By rigid deformations, we mean additional contributions that arise due to the worldsheet isometries or are related to displacements of a hypersurface via isometries of the embedding spacetime which preserve its extrinsic curvature. [^1] We discuss possible physical interpretations of the rigid deformations especially in Sec. 5. In particular, we show that the Nambu-Goto equation is recovered in the tensionless limit in absence of the rigid deformations.
@@ 2 Setup and statement of results
2.1 The question and why $D=3$ is special
Consider a junction $\Sigma$ constituted by a co-dimension one brane with tension $T_{0}$ embedded in a $D$-dimensional manifold $\mathcal{M}$. The full manifold $\mathcal{M}$ involves the union of two $D$-dimensional spacetimes $\mathcal{M}_{L}$ and $\mathcal{M}_{R}$ which are glued at the brane junction $\Sigma$. The Einstein-Hilbert action with a cosmological constant with the brane source is
$S=\frac{1}{16\pi G_{N}}\int_{\mathcal{M}}{\rm d}^{D}x\sqrt{-g}(R+2\Lambda)+T_{0}\int_{\Sigma}{\rm d}^{D-1}y\sqrt{-\gamma}+{\rm GHY}\,\,{\rm terms},$ (1)
where $g$ is the bulk metric, $\gamma$ is the induced metric on $\Sigma$ and GHY terms stand for the Gibbons-Hawking-York boundary terms. [^3] There are two GHY terms associated to $\Sigma$ as it is the part of the boundary of both $\mathcal{M}_{L}$ and $\mathcal{M}_{R}$. If there are additional boundaries, then there are more GHY terms. However, the latter will not be relevant for the current discussion. It is to be noted that the bulk metric $g$ is the only physical field in this action. The embedding of $\Sigma$ should be determined from the junction conditions obtained from extremizing this action with respect to variations of the bulk metric at the junction, and not from separate equations of motion. Therefore, the limit $T_{0}\rightarrow 0$ need not correspond to the probe brane in an Einstein manifold.
The natural question is whether we can identify a generic solution of the gravitational theory described by the action (1) with solutions of the worldvolume extremization equations of a co-dimension one brane in $\mathcal{M}$. We can readily argue this is not possible for $D\geq 4$, and that the case $D=3$ is special, as follows.
Instead of using continuous bulk coordinates in which the normal and the tangents of $\Sigma$ are continuous across the junction, it is convenient for our purposes to adopt the methodology of Bachas:2020yxv (but without restricting to only linear perturbations about a simple solution and assuming any specific boundary conditions). Following Bachas:2020yxv , we adopt coordinates $(t_{L},x_{L},\vec{z}_{L})$ for $\mathcal{M}_{L}$ and $(t_{R},x_{R},\vec{z}_{R})$ for $\mathcal{M}_{R}$, where $t_{L,R}$ are the time coordinates, while $x_{L,R}$ and the $\vec{z}_{L,R}$ with $D-2$ components are the $D-1$ spatial coordinates in the respective halves. The hypersurface $\Sigma$ has two images in $\mathcal{M}_{L}$ and $\mathcal{M}_{R}$, which are $\Sigma_{L}$ and $\Sigma_{R}$ respectively, and should be glued via solving the junction conditions. The hypersurface $\Sigma_{L}$ in $\mathcal{M}_{R}$ is given by
$x_{L}=f_{L}(t_{L},\vec{z}_{L}),$ (2)
Similarly, the hypersurface $\Sigma_{R}$ in $\mathcal{M}_{R}$ is given by
$x_{R}=f_{R}(t_{R},\vec{z}_{R}).$ (3)
Furthermore, $\mathcal{M}_{L}$ is the set of points $x_{L}\leq f_{L}(t_{L},\vec{z}_{L})$ while $\mathcal{M}_{R}$ is the set of points $x_{R}\geq f_{R}(t_{R},\vec{z}_{R})$.
A point $P_{L}$ on $\Sigma_{L}$ is identified with another point $P_{R}$ on $\Sigma_{R}$ if these points carry the same brane coordinate labels $(\tau,\vec{\sigma})$. However, we also need to fix the diffeomorphisms on the brane. It is convenient to fix the diffeomorphisms on the brane via these gauge conditions:
$\tau=\frac{1}{2}(t_{L}(P_{L})+t_{R}(P_{R})),\quad\vec{\sigma}=\frac{1}{2}(\vec{z}_{L}(P_{L})+\vec{z}_{R}(P_{R})).$ (4)
Therefore, the parametric forms of $\Sigma_{L}$ and $\Sigma_{R}$ are
$t_{L}(\tau,\vec{\sigma})=\tau-\tau_{a}(\tau,\vec{\sigma}),\quad\vec{z}_{L}(\tau,\vec{\sigma})=\vec{\sigma}-\vec{\sigma}_{a}(\tau,\vec{\sigma}),\quad x_{L}(\tau,\vec{\sigma})=f_{L}\left(t_{L}(\tau,\vec{\sigma}),\vec{z}_{L}(\tau,\vec{\sigma})\right),$
$t_{R}(\tau,\vec{\sigma})=\tau+\tau_{a}(\tau,\vec{\sigma}),\quad\vec{z}_{R}(\tau,\vec{\sigma})=\vec{\sigma}+\vec{\sigma}_{a}(\tau,\vec{\sigma}),\quad x_{R}(\tau,\vec{\sigma})=f_{R}\left(t_{R}(\tau,\vec{\sigma}),\vec{z}_{R}(\tau,\vec{\sigma})\right),$ (5)
where $\tau_{a}=(1/2)(t_{R}-t_{L})$ and $\vec{\sigma}_{a}=(1/2)(\vec{z}_{R}-\vec{z}_{L})$ are the relative time and spatial shifts which are not fixed by the choice of gauge and should be determined via the junction conditions. In total, the junction conditions need to determine $D+1$ functions of $\tau$ and $\vec{\sigma}$, namely the $D-1$ relative coordinates $\tau_{a}$ and $\vec{\sigma}_{a}$, and $f_{L}$ and $f_{R}$ which give the images of the embedding of the brane. Solving the $D+1$ variables we can determine the geometric embedding of the brane in the bulk spacetime completely.
Explicitly by extremizing the action (1), we obtain Einstein’s equations
$R_{MN}-\frac{1}{2}Rg_{MN}+\Lambda g_{MN}=0,$ (6)
away from $\Sigma$, i.e. away from $\Sigma_{L}$ and $\Sigma_{R}$ in $\mathcal{M}_{L}$ and $\mathcal{M}_{R}$, respectively. Above, $M$ and $N$ stand for the $D$ bulk indices. The extremization on the junction gives Israel:1966rt
$\gamma_{\mu\nu}|_{\Sigma_{R}}-\gamma_{\mu\nu}|_{\Sigma_{L}}$ $=$ $0,$ (7)
$(K_{\mu\nu}-K\gamma_{\mu\nu})_{\Sigma_{R}}-(K_{\mu\nu}-K\gamma_{\mu\nu})_{\Sigma_{L}}$ $=$ $8\pi G_{N}T_{0}\gamma_{\mu\nu},,$ (8)
where $\mu$ and $\nu$ are the $D-1$ brane indices, and $K_{\mu\nu}$ is the (respective hypersurface’s) extrinsic curvature. The above are called the Israel junction conditions Israel:1966rt . The combination $K_{\mu\nu}-K\gamma_{\mu\nu}$ is also called the Brown-York stress tensor Brown:1992br .
The first set of junction conditions (7) involving the continuity of the induced metric give $D(D-1)/2$ equations. Although the second set of junction conditions (8) involving the discontinuity of the Brown-York stress tensor give $D(D-1)/2$ equations as well, $D-1$ of them are redundant as the Brown-York stress tensor of any hypersurface is identically conserved (in the background given by the induced metric) on an Einstein manifold by virtue of the Gauss-Codazzi equations. Therefore, the junction conditions give in total
$D(D-1)-(D-1)=(D-1)^{2}$
equations. However, we have already seen that the embedding of the junction is determined by $D+1$ variables. We readily observe that for $D\geq 3$,
$(D-1)^{2}\geq D+1,$
with this inequality saturated only when $D=3$. Therefore, for $D\geq 4$, there are more equations that are obtained from the junction conditions than the number of variables which determine the geometric embedding of the brane. This argument implies that we cannot recover generic solutions of the worldvolume extremization equations of the brane from the junction conditions in the tensionless limit for $D\geq 4$.
The case $D=3$ is special as the junction conditions give 4 equations that can fully determine the 4 variables which in turn completely specify the geometric embedding of the brane. In this paper, we will show that the naive counting argument is indeed correct and we can identify any solution of the gravitational equations as the backreaction of a string which satisfies the Nambu-Goto equations in the probe limit ($T_{0}\rightarrow 0$) up to a finite number of rigid deformation parameters in this case. We will describe our setup and summarize our results precisely in the following subsection. [^4] The case $D=2$ is also special. It can be shown that in this case, the general solutions involve gluing two copies of flat spaces ($\Lambda=0$), or two AdS2 ($\Lambda<0$), or two dS2 ($\Lambda>0$) spacetimes with identical metrics, but with an isometry performed on one of the two copies. The junction follows the geodesic equation and this geodesic is invariant under a one parameter family of isometries. The isometry which is operated on one of the two bulk copies of flat space, AdS2 or dS2 is within this one parameter family, and is determined by the conserved charge of the geodesic. (For every isometry, a geodesic has a conserved charge.)
2.2 Setup and results
In this paper, we study the solutions of the action (1) in $D=3$. However we also further specialize to the case of gluing two identical copies of an Einstein spacetime satisfying the equations (6) for the simplicity of understanding how the Nambu-Goto equation arises from the junction conditions. For the sake of clarity, we describe our setup from scratch instead of falling back upon the general discussion in the previous subsection although many aspects of the discussion are going to be repetitive.
@@ 2.2 Setup and results
Consider two copies of a $2+1$-dimensional Einstein manifold $\mathcal{M}$, which is locally flat or an AdS space. We label the copies of $\mathcal{M}$ as $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$. Let $\Sigma_{1}$ and $\Sigma_{2}$ be two $1+1$-dimensional hypersurfaces embedded in $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$, respectively, and splitting each of these spacetimes into two halves. We will study solutions of Einstein’s equations obtained by gluing one of the halves of $\mathcal{M}_{1}$ with one of the halves of $\mathcal{M}_{2}$ by identifying points in $\Sigma_{1}$ and $\Sigma_{2}$ such that the induced metrics and extrinsic curvatures satisfy the Israel junction conditions Israel:1966rt . As mentioned before, our methodology is similar to that adopted in Bachas:2020yxv , except that allow more general boundary conditions and also not restrict ourselves to linear perturbations about a simple solution. See Fig. 1 for an illustration of the setup in AdS.
@@
$t_{1}(\tau,\sigma)=\tau-\tau_{a}(\tau,\sigma),\quad z_{1}(\tau,\sigma)=\sigma-\sigma_{a}(\tau,\sigma),\quad x_{1}(\tau,\sigma)=f_{1}\left(t_{1}(\tau,\sigma),z_{1}(\tau,\sigma)\right),$
$t_{2}(\tau,\sigma)=\tau+\tau_{a}(\tau,\sigma),\quad z_{2}(\tau,\sigma)=\sigma+\sigma_{a}(\tau,\sigma),\quad x_{2}(\tau,\sigma)=f_{2}\left(t_{2}(\tau,\sigma),z_{2}(\tau,\sigma)\right).$ (12)
@@
$x_{s}(\tau,\sigma)=\frac{1}{2}\left(f_{1}\left(t_{1}(\tau,\sigma),z_{1}(\tau,\sigma)\right)+f_{2}\left(t_{2}(\tau,\sigma),z_{2}(\tau,\sigma)\right)\right),$
$x_{a}(\tau,\sigma)=\frac{1}{2}\left(f_{2}\left(t_{2}(\tau,\sigma),z_{2}(\tau,\sigma)\right)-f_{1}\left(t_{1}(\tau,\sigma),z_{1}(\tau,\sigma)\right)\right),$ (13)
@@
Henceforth, we will refer to $\lambda$ as the tension. The two junction conditions give six equations. However, the Brown-York stress tensor is identically conserved on any hypersurface of an Einstein manifold by virtue of the Gauss-Codazzi equations and this guarantees that the traceless part of the Brown-York tensor is continuous at the junction even when $\lambda\neq 0$. Therefore, the gluing conditions (15) give only one independent equation which together with (14) amount to four equations for determining the four variables, namely $\tau_{a}$, $\sigma_{a}$, $f_{1}$ and $f_{2}$.
We will see that a solution of the junction conditions depends on the tension $\lambda$ and a finite number of integration constants $\delta_{i}$, that we call rigid deformations. The latter are associated with worldsheet isometries and rigid displacements of the hypersurface via isometries of the embedding spacetime which do not change its extrinsic curvature.
@@ 3.1 Solving the junction conditions perturbatively
It is convenient to use the average and relative coordinates $x_{s}$ and $x_{a}$ as defined in (2.2). Obviously, the coefficients of their perturbative expansions are
@@
where $\dot{\phantom{a}}$ and $\phantom{a}{}^{\prime}$ imply partial derivatives w.r.t. $\tau$ and $\sigma$, respectively. At this order, the junction conditions (15) give
@@
It is easy to see that $\alpha_{k}$, $\beta_{k}$ and $\gamma_{k}$ are rigid displacement parameters that preserve the extrinsic curvature of a hypersurface. Explicitly, these correspond to infinitesimal isometries of the embedding flat space which do not keep the hypersurface $x=0$ (where $\Sigma_{1}$ and $\Sigma_{1}$ coincide at the zeroth order) invariant; $\alpha_{k}$ is a transverse spatial displacement, $\beta_{k}$ is an infinitesimal rotation and $\gamma_{k}$ is an infinitesimal transverse boost. The other three isometries, namely time translation, longitudinal spatial displacement and the longitudinal boost keep $x=0$ invariant.
It will turn out that the higher orders in the expansion will not introduce new parameters, so that in total $\alpha_{h}$, $\beta_{h}$, $\gamma_{h}$, $\alpha_{k}$, $\beta_{k}$ and $\gamma_{k}$ give six rigid parameters which together with the initial conditions for $x_{s}$ uniquely specify a solution of the junction conditions to all orders. By assumption, all these six rigid parameters are small, i.e. $\mathcal{O}(\varepsilon)$. Although these infinitesimal rigid parameters are related to worldsheet and spacetime isometries, at higher orders they generically affect the full non-linear solution non-trivially. Furthermore, these parameters should be determined by appropriate boundary conditions depending on the physical context, and therefore can generally be physical parameters. We will discuss boundary conditions explicitly in the context of junctions of AdS in the next section.
@@ 4.1 Solving the junction conditions perturbatively
One crucial point is that we do not impose any boundary conditions on $\tau_{a}$ and $\sigma_{a}$. In fact, these will be given by two rigid parameters related to worldsheet isometries at the leading order. [^6] This is in contrast to the setup of Bachas:2020yxv , where they imposed Dirichlet boundary conditions on all the variables and found a unique solution for each given frequency in pure AdS. The diffeomorphisms related to these two parameters preserve the worldsheet boundary at $\sigma=\infty$ but induce relative time-reparametrization at the junction since $\tau_{a}$ does not vanish at the boundary of AdS.
Unlike the case of flat space, our perturbative expansion will be valid near the boundary although it will be unreliable near the horizon.
@@
[…]
Specifically, imposing the Dirichlet boundary condition on both $x_{1}$ and $x_{2}$, leads to the boundary conditions (62) for $x_{a,1}$ implying that it should vanish at $\sigma\rightarrow\infty$. This can be satisfied if
@@ 5 Discussion
Our general solutions which correspond to the Nambu-Goto solutions up to the rigid deformations should be interpreted as general state-dependent interfaces in the dual conformal field theory. Note that the Nambu-Goto equation has no intrinsic dimensionful parameters, and the quasinormal mode spectrum that characterize the solutions is determined by the mass of the background spacetime, i.e. the dual state. When we impose the Dirichlet boundary condition identically on both the hypersurfaces (i.e. for both $x_{s}$ and $x_{a}$), the solutions should holographically represent dynamical interfaces between two copies of the same state in the dual CFT, particularly two thermal states with the same temperature since the BTZ black hole is dual to a thermal state. No transfer of energy and momentum occur through the interfaces. Nevertheless, the interfaces have their own dynamics, which are characterized by the solutions of the Nambu-Goto equations and the rigid parameters of the bulk solution. All these parameters and the solutions of the Nambu-Goto equation together with the tension determine the correlation functions of the full system. In this aspect, it could be important to recall that two sides of the interface necessarily have a relative time-reparametrization (given by the asymptotic limit of $\tau_{a}$) which can decode the non-trivial Nambu-Goto solution in the bulk. This feature is not present in a usual defect. In this context, it is important to understand such interfaces in the vacuum which correspond to bulk junctions in pure AdS3.
The holographic interpretation of our solutions would be important to understand whether the rigid parameters are related to the physical properties of the interface. The appropriate boundary conditions for the junction corresponding to the dual interface set the values of some of the rigid physical parameters (as we have seen in the context of the Dirichlet boundary condition in AdS). The remaining rigid parameters may actually characterize physical properties of the interface. Although the latter can be absorbed via relative isometries at the leading order, they might correspond to physically inequivalent solutions (as in the case of improper diffeomorphisms in AdS/CFT). In the context of AdS, these issues can also be addressed by the computation of holographic entanglement and Renyi entropies, which can directly reveal the full bi-partite state in the CFT glued at the interface, and also clarify whether the tensionless limit leads to a trivial interface.
@@
A related question is whether we can go beyond a perturbative approach and formulate a way to solve the gravitational equations with the junction conditions via numerical relativity setting necessary initial conditions and also boundary conditions. In fact, a fundamental question in this context is which types of initial data would lead to non-singular spacetimes such as those not having closed time-like curves. Such an initial value formulation should lead to the most general constructions of gravitational solutions involving junctions which can be interpreted as novel interfaces holographically.